If h is plankcs's constant,the momentum of a photon of wavelength 1anstrom in?
1
Expert's answer
2019-03-13T11:54:24-0400

Solution:

"p=h\/\u03bb"


p = (6.62 × 10-34 J/s) / (1 × 10-10 m) = 6.62×10-24  Jm/s

Answer: 6.62×10-24  Jm/s